One way the U.S. Environmental Protection Agency (EPA) tests for chloride contaminants in water is by titrating a sample of silver nitrate solution. Any chloride anions in solution will combine with the silver cations to produce bright white silver chloride precipitate. Suppose an EPA chemist tests a 250.mL sample of groundwater known to be contaminated with iron(II) chloride, which would react with silver nitrate solution like this: FeCl2(aq) + 2AgNO3(aq) → 2AgCl(s) + FeNO32(aq) The chemist adds 54.0mM silver nitrate solution to the sample until silver chloride stops forming. She then washes, dries, and weighs the precipitate. She finds she has collected 3.4mg of silver chloride. Calculate the concentration of iron(II) chloride contaminant in the original groundwater sample. Be sure your answer has the correct number of significant digits.
Answer #1


FeCl2(aq) + 2AgNO3(aq) ----> 2AgCl(s) + FeNO32(aq)

mass of precipitate(AgCl) formed = 3.4 mg

No of mole of AgCl formed = 3.4/143.32 = 0.0237 mmole

1 mole AgCl = 1 mole Cl-

No of mole of Cl- present in sample = 0.0237 mmole

1 mole FeCl2 = 2 mole Cl-

No of mole of FeCl2 present in sample = 0.0237*(1/2) = 0.0119 mmole

concentration of FeCl2 in sample = n/v

                                 = 0.0119/250

                                 = 4.76*10^-5 M